OpenStudy (ria23):

>n< Halp. Expand from vertex to standard. y-2=-3(x+4)^2 >.< I expanded it, and ended up cancelling my x's... I ended up with y=-3x^2+14..... Not only is that not standard form, but I know that my answer is supposed to be y=-3x^2-24x-46 >n<

OpenStudy (xapproachesinfinity):

well just expand what's the problem here?

OpenStudy (ria23):

The problem is y-2=-3(x+4)^2 I expanded it to y-2=-3(x+2)(x+2) And it cancelled out my x...

OpenStudy (xapproachesinfinity):

oh i see so missed something when you expanded why x term disappear? shouldn't be that way

OpenStudy (xapproachesinfinity):

first of all that's not true (x+4)^2=(x+4)(x+4)

OpenStudy (xapproachesinfinity):

but even so i don't think you x's will cancel even with that mistake but the answer will turn to be wrong since you use a wrong expansion

OpenStudy (ria23):

There's a...... Different expansion? O.o

OpenStudy (xapproachesinfinity):

No there isn'tNo there isn't what you did is wrong (a+b)^2=(a+b)(a+b) see here i haven't changed my a and b? you problem is that you are thinking about a^2-b^2=(a+b)(a-b) although there totally different but you thought about something like this x^2-9=(x+3)(x-3) what you did is wrong (a+b)^2=(a+b)(a+b) see here i haven't changed my a and b? you problem is that you are thinking about a^2-b^2=(a+b)(a-b) although there totally different but you thought about something like this x^2-9=(x+3)(x-3)

OpenStudy (xapproachesinfinity):

hehe OP is going crazy lol

OpenStudy (xapproachesinfinity):

Sorry for the repeatition open study got stuck

OpenStudy (ria23):

That's alright. So, Like yhu said, if I expand, it should look like y-2=-3(x+4)(x+4) Right?

OpenStudy (xapproachesinfinity):

correct now distribute that parenthesis

OpenStudy (ria23):

\[y-2=(-3-12)(x+4)\] ... Right?

OpenStudy (ria23):

Oops I forgot my x... That should say (-3x-12)*

OpenStudy (xapproachesinfinity):

yeah! continue

OpenStudy (ria23):

\[y-2=-3x ^{2}-24x-48\] >O<; Oh my god. Thank yhu. ^u^